Historical status reports for 2007

16:40    Due to a fibre optic cable cut in the USA, international connectivity
	        is somewhat patchy at the moment. Some sites are unaffected, some are
         slow and some are unreachable. The problem is widespread, and being
         worked on as a matter of urgency. No ETR at this stage.
10:35    There will be a brief (2 minute) interruption to all authentication
         of all ADSL services while equipment is upgraded. (Update: completed
         without incident)
01:30    Loss of all NSW adsl services. Wholesaler investigating reason.
         Services restored 8:05am approx but still no reason.
21:10    Emergency shutdown of the Sydney LNS (affecting all our DSL customers
         and some dial-up users) required to replace faulty flash memory module.
         Anticipated outage is under 5 minutes.
07:30    Loss of internet/transit for adsl and national roaming customers due
         to the failure of interface card in upstream providers router.
         Customers will have had access to mail, our network and a few other
         sites who peer in Sydney.
         Service restored approx 08:58
07:00    Between 07:00 and 07:05 this morning, we lost virtually all our Victorian
         ADSL services. Notice from Optus is as follows:
           Service Impact:  Services in VIC
           Expected Duration:  Unknown
           Work Description:  A power issue in the Melbourne POP is causing
                interruption to VIC services.  There is no current ETA.  
                Updates will be provided as they become available.
11:45    Secondary nameserver has had a HDD failure.
         Building / restoring replacement system.
12:03    We are experiencing problems with national dialup for subscribers in
         MtBeauty, Bright and Falls Creek. 
         (Update: 14:50 - Optus advise there are widespread problems in Victoria
         No ETA but they are working on the issue)
02:00    Telstra have advised that "Due to Software Upgrade in the exchange"
P439951/BN there will be a 5-minute interruption to all our Albury and Mt Beauty
         dial-in numbers. This interruption should not affect those using our
         National number.
11:55    Power works are completed, all equipment returned to the normal
         supplies, no interruption or disruption to service throught the

06:00    Due to major electrical works being done in the Griffith Road area,
	 we have been advised there will be a power outage for approx 6 hours.
         We have made arrangements to supply our office on alternate power, and
         do not expect this interruption to affect our services in any way.
         (A test run has been completed from 11:10 to 14:15 20/Jun/2007)
14:00	 Old faithful orac, our longest-running server suffered a fan failure
	 in its power supply, cooking the supply. Orac was still running but
	 has been badly stressed due to the lack of airflow. After 12 years
	 and 3 days of faithful service with a total downtime of under an hour
	 (mostly for relocation or additional serial ports), and a last "uptime"
	 of 1478 days, we've turned orac off for a well earned rest.

05:00    We have been advised by our wholesaler of a planned interruption to all
         ADSL services for up to 5 minutes while they increase capacity.
14:35    Complete loss of connectivity to SDH Access point.
	 Affected services: skycam, weatherstation, lightning tracking, some
         phone services and sysadmin. After an uptime of 991 days, the wireless
         card in the AP locked up, requiring a hard reset and reconfig. Odd.
         Services restored 15:20
12:55    Power returned, site operational. Unsure of power stability.

12:29    Widespread powerfail affecting large parts of northeast victoria.
	 As a result, our Mount Beauty site has been running on UPS, but has now
         expired, taking the POP off air. Electricity authority advises they are
         hoping to have service restored about 14:30
	 Wangaratta POP decommissioned. All users should by now have completed
	 the migration process mailed or given verbally to all affected users.

         Modem card on the access server in Corryong shutdown due to overtemp.
         The site has been temporarily unattended, aircon was off and rack temp
         got excessive to the point the card shutdown on overtemperature.
         Aircon sorted, temp brought back to reasonable levels and all ok.